What is the point that is common to the graph on both functions?
Y1 = 1/2 X+3
Y2 = -X-3

Answer by stanbon(75887)

What is the point that is common to the graph on both functions?
Y1 = 1/2 X+3
Y2 = -X-3
----
The y-values would have to be the same.
So (1/2)x + 3 = -x-3
(3/2)x = -6
x = (2/3)(-6)
x = -4
---
Solve for "y":
y = -x-3
y = 4-3 = 1
----
Common Point (-4,1)
========================
Cheers,
Stan H.
========================